# BoViLA: Bootstrapping Video-Language Alignment via LLM-Based Self-Questioning and Answering

### v1

@@ Introduction
However, directly applying vanilla EDL to LLMs would fail. This is because vanilla EDL commonly applies non-negative activation functions like ReLU(Sensoy, Kaplan, and Kandemir 2018) or Softplus(Amini et al. 2020) on logits to ensure the non-negativity of evidence. This causes information loss and negative effects(Ye et al. 2024; Meinert, Gawlikowski, and Lavin 2023; Wu et al. 2024b), especially for finetuning pretrained models with a large number of categories. To overcome this and successfully apply EDL to LLMs, we propose decoupling the direction and magnitude of the evidence vector $\bm{e}=(e_{1},e_{2},\cdots,e_{K})$ to mitigate information loss. To be detailed, assume the logits output by the model are $\bm{z}=(z_{1},z_{2},\cdots,z_{K})$, vanilla EDL determines $\bm{\alpha}$ through the following transformation:
@@
$=-\sum_{j=1}^{N_{a}}\int\log p_{j{a_{j}}}\frac{1}{B(\bm{\alpha}_{j})}\prod_{i=1}^{K}p_{ji}^{\alpha_{ji}-1}d\bm{p}_{j},$ (17)
@@
We also apply the regularization loss $\mathcal{L}_{\mathrm{reg}}^{\mathrm{edl}}$ mentioned in (Sensoy, Kaplan, and Kandemir 2018). After training, EDL uncertainty can be estimated as:
@@
where $S_{i}$ represents total strength corresponding to $p(\cdot|\bm{v},\bm{q},a_{<i})$. This uncertainty is used to filter self-generated questions of low-quality by simply controling the weight of loss $\mathcal{L}_{\mathrm{v\overline{q}a}}$ with $1-u$. Finally, we train BoViLA with the following total loss:
@@
Experimental Setup
Implementation Details.
We conduct all training with 8 × 80GB A800 GPUs for 10 epochs. For all the datasets, we use VIT-L/14 as the visual encoder to extract 10 frame features for each video and use LLaMA(7B) as our large language model. Regarding evaluation metrics, we use the accuracy of choosing the right answer and test on the validation split. We provide the detailed prompt template of BoViLA in Table 1. Please refer to Appendix for more training details.
Table 1: Input Prompt of Questioner and Answerer.
Baselines & Benchmarks.
We compare our framework with some state-of-the-art (SOTA) baselines, especially that are LLM-based such as BLIP-2(Li et al. 2023), SeViLA(Yu et al. 2024) and LLaMA-VQA(Ko et al. 2023), on 5 challenging multi-choice VideoQA benchmarks: 1) TVQA(Lei et al. 2018), requireing answer questions based on video, dialogues and scenes, featuring 152,545 QA pairs from 21,793 video clips extracted from popular TV shows. 2) STAR(Wu et al. 2024a), which is designed for spatio-temporal and relational reasoning, containing 22,670 QA pairs based on 12,672 video clips. 3) DramaQA(Choi et al. 2021), which is tailored for emotional and social reasoning, featuring 16,191 QA pairs derived from 23,239 video clips. 4) VLEP(Lei et al. 2020), which focuses on predicting future events based on video and dialogues, consisting of 28,726 QA pairs from 10,000 video clips. 5) How2QA(Li et al. 2020), which is designed for instructional video comprehension, containing 46,467 QA pairs derived from 23,228 video clips.
Main Results
Models Language Model # trainable params STAR DramaQA VLEP TVQA* How2QA Int. Seq. Pre. Fea. Tot. Tot. Tot. Tot. Tot. FrozenBiLM (Yang et al. 2022) DeBERTa 30M - - - - - - - 57.5 86.7 MERLOT (Zha et al. 2019) RoBERTa 223M - - - - - 81.4 68.4 - - SPCRL (Kim et al. 2021) BERT - - - - - - 81.0 - - - AIO (Wang et al. 2023a) - 110M 47.5 50.8 47.8 44.1 47.5 - - - - ATP (Buch et al. 2022) CLIP - 50.6 52.9 49.4 40.6 48.4 - - - - MIST (Gao et al. 2023) - - 55.6 54.2 54.2 44.5 53.9 - - - - InternVideo (Wang et al. 2022b) CLIP 1.3B 62.7 65.6 54.9 51.9 58.7 - 63.9 57.2 79.0 LLaMa-VQA LLaMA 4.5M 66.2 67.9 57.2 52.7 65.4 84.1 71.0 70.4 - BLIP-2 Flan-T5 432M 52.3 54.8 49.0 51.2 51.8 - 67.0 54.5 82.2 SeViLA Flan-T5 216M 63.7 70.4 63.1 62.4 64.9 - 68.9 61.6 83.6 VLAP Flan-T5 188M 70.0 70.4 65.9 62.2 67.1 - 69.6 63.4 83.9 BoViLA (Ours) LLaMA 4.5M 66.9 68.0 62.0 57.2 66.4 85.2 71.2 71.6 89.4
@@
As shown in Figure 3, the self-questioning and answering process can easily cause ’ information leakage’, where the questioner creates degenerate questions, which are meaningless and cheatingly contain target answers. Our regularization method and EDL-based filter really help solve this problem and allow the model to generate high-quality questions to boost answers. Furthermore, we provide comprehensive ablation studies on STAR validation set in Table 3 to verify the effectiveness of our method. The results clearly demonstrate: 1) Only learning from regularization about questioning allows the model to achieve a 2.2% performance boost. 2) Further learning by answering these self-generated questions contributes an additional 0.7% improvement. 3) If the answerer is allowed to backpropagate gradients to the questioner, i.e. providing feedback on question quality, performance can increase by another 0.5%. 4) Moreover, applying an EDL-based filter to progressively eliminate potential junk questions that could negatively impact the model can further enhance performance by 0.9%.
Figure 3: Examples of self-questioning. Seed Question refers to the questions labeled in the dataset, and Answer represents corresponding answers along with respective options. Degenerate Question refers to the junk questions generated by the model when regularization and EDL-based filter are not used. Self-Generated Question 1 and 2 represent high-quality questions generated by the model equipped with regularization and EDL-based filter. These questions are semantically consistent with the video and leverage the internal knowledge of the LLMs to provide semantically similar but different questions. The uncertainty is estimated by the EDL head. In these two examples, it is evident that the question quality is negatively correlated with the uncertainty. Additionally, it is worth mentioning that since we used the gumbel-softmax function with some randomness in sampling self-generated questions, different questions are generated for the same sample during each training epoch.
Linear Softmax STAR Int. Seq. Pre. Fea. Avg. - - - - - - - - 24.7 23.1 22.9 21.6 23.5 66.9 68.0 62.0 57.2 66.4
@@
We also conduct an ablation study on our improved EDL method. As shown in Table 4, without our proposed linear projection to calculate the total strength yet simply sum it up, the model’s training will break down. This is due to the large number of classes in the vocabulary (even though most words receive very small logits), and summing directly without considering sparsity can easily lead to numerical instability. Moreover, if we do not use our proposed Softmax-based method to decouple the computation of evidence direction, the significant information loss will also cause the model to fail to converge.
More Discussion
Our proposed EDL-based filter is built on the assumption that “EDL-estimated uncertainty is able to measure the quality of self-generated questions and, is approximately negatively correlated with it”. Here, we further explore and validate this assumption. We use $\mathcal{L}_{\mathrm{v\overline{q}a}}$ and $\mathcal{L}_{\mathrm{reg}}$ as approximate measures of the quality score for self-generated questions. $\mathcal{L}_{\mathrm{v\overline{q}a}}$ represents the similarity between the self-generated question $\overline{q}$ and the seed question $q$ in terms of KL divergence, where a high $\mathcal{L}_{\mathrm{v\overline{q}a}}$ indicates a completely uncontrolled self-generated question. $\mathcal{L}_{\mathrm{reg}}$ represents the likelihood that the answerer successfully predicts the target answer from the self-generated question, where a high $\mathcal{L}_{\mathrm{reg}}$ is likely to indicates that there is no meaningful semantics in the self-generated question. $\mathcal{L}_{\mathrm{v\overline{q}a}}$ and $\mathcal{L}_{\mathrm{reg}}$ complement and reinforce each other, so we use them together to represent the quality of self-generated questions. As shown in Figure 4, both $\mathcal{L}_{\mathrm{v\overline{q}a}}$ and $\mathcal{L}_{\mathrm{reg}}$ demonstrate the pattern that ”the lower the quality of the self-generated question, the higher the uncertainty”. This insight further provides a solid explanation for the effectiveness of our proposed EDL-based filter.
@@
Conclusion, Limitation and Future Work
In this paper, we present BoViLA, a pioneering framework that enhances video-language alignment through self-questioning and answering. BoViLA utilizes a unique bootstrapping approach where the model alternates between the roles of questioner and answerer, enabling it to generate and answer self-created questions, thereby deepening modality alignment without reliance on extra annotated data by fully leverage thr rich information within videos and internal knowledge in LLMs. Our framework also improve the vanilla EDL method and incorporates it to assess and filter the quality of self-generated questions. BoViLA demonstrates superior performance across five VideoQA benchmarks, outperforming current state-of-the-art methods with only a few trainable parameters.
Despite its success, the framework faces limitations in a constrained question generation space because our questioner always generates new questions by autoregressively predicting the next token based on the context of the seed question. Future work will focus on exploring how to sample more freely from the joint distribution of question samples during training, in order to generate more diverse self-generated questions.
@@ Appendix B Details of EDL Loss
$\mathcal{L}_{\mathrm{vqa}}^{\mathrm{edl}}=\sum_{j=1}^{N_{a}}E_{\mathrm{Dir}}[-\log p_{j{a_{j}}}]=\sum_{j=1}^{N_{a}}\int-\log p_{j{a_{j}}}\frac{1}{B(\bm{\alpha}_{j})}\prod_{i=1}^{K}p_{ji}^{\alpha_{ji}-1}d\bm{p}_{j}.$ (A)
@@
$\mathcal{L}_{\mathrm{reg}}^{\mathrm{edl}}$ $=KL[Dir(\bm{p}|\bm{\alpha})\|Dir(\bm{p}|\langle 1,\ldots,1\rangle)]$ (D)
$=E_{Dir(\bm{p}|\bm{\alpha})}\left[\log\frac{Dir(\bm{p}|\bm{\alpha})}{Dir(\bm{p}|\langle 1,\ldots,1\rangle))}\right]$ (E)
$=E_{Dir(\bm{p}|\bm{\alpha})}\left[\log Dir(\bm{p}|\bm{\alpha})-\log Dir(\bm{p}|\langle 1,\ldots,1\rangle)\right)]$ (F)
$=E_{Dir(\bm{p}|\bm{\alpha})}\left[-\log B(\bm{\alpha})+\sum_{k=1}^{K}(\alpha_{k}-1)\log p_{k}\right]+\log B(\langle 1,\ldots,1\rangle)$ (G)
$=\log\frac{B(\langle 1,\ldots,1\rangle)}{B(\bm{\alpha})}+E_{Dir(\bm{p}|\bm{\alpha})}\left[\sum_{k=1}^{K}(\alpha_{k}-1)\log p_{k}\right]$ (H)
@@ Appendix C Validity of EDL-Estimated Uncertainty
Visualization of Uncertainty Distribution.
@@
Appendix D Case Study about Different Forms of question
We present here the model’s varying responses when confronted with semantically consistent but differently formatted questions, as shown in Figure C, which demonstrates the necessity of our approach.
Figure C: Limitations of the model in facing different forms of questions.

### v2

@@ 3.3 EDL-based Filter
However, directly applying vanilla EDL to LLMs would fail. This is because vanilla EDL commonly applies non-negative activation functions like ReLU (Sensoy et al., 2018) or Softplus (Amini et al., 2020) on logits to ensure the non-negativity of evidence. This causes information loss and negative effects (Ye et al., 2024; Meinert et al., 2023; Wu et al., 2024b), especially for finetuning pretrained models with a large number of categories. To overcome this and successfully apply EDL to LLMs, we propose evidence decoupling to decouple the direction and magnitude of the evidence vector $\bm{e}=(e_{1},e_{2},\cdots,e_{K})$ to mitigate information loss. To be detailed, assume the logits output by the model are $\bm{z}=(z_{1},z_{2},\cdots,z_{K})$, vanilla EDL determines $\bm{\alpha}$ through the following transformation:
@@
$=-\sum_{j=1}^{N_{a}}\int\log p_{j{a_{j}}}\frac{1}{B(\boldsymbol{\alpha}_{j})}\prod_{i=1}^{K}p_{ji}^{\alpha_{ji}-1}d\bm{p}_{j},$ (17)
@@
We also apply the regularization loss $\mathcal{L}_{\mathrm{reg}}^{\mathrm{edl}}$ mentioned in (Sensoy et al., 2018). After training, EDL uncertainty can be estimated as:
@@
where $S_{i}$ represents total strength corresponding to $p(\cdot|\bm{v},\bm{q},a_{<i})$.
To the best of our knowledge, our work is the first to successfully apply EDL to LLMs, offering a more convenient method for measuring the uncertainty of LLMs. We conduct more detailed ablation studies and analysis on the effectiveness of our EDL-based filter and the evidence decoupling technique in Section 4.4 and the supplementary materials.
This uncertainty is then used to filter self-generated questions of low-quality by simply controling the weight of loss $\mathcal{L}_{\mathrm{v\overline{q}a}}$ with $1-u$. Finally, we train BoViLA with the following total loss:
@@ 4 Experiment
4.1 Experimental Setup
4.1.1 Implementation Details.
We conduct all training with 8 × 80GB A800 GPUs for 10 epochs. For all the datasets, we use VIT-L/14 as the visual encoder to extract 10 frame features for each video and use LLaMA(7B) as our large language model. Regarding evaluation metrics, we use the accuracy of choosing the right answer and test on the validation split. More details are provided in Appendix A.
4.1.2 Baselines & Benchmarks.
We compare our framework with some state-of-the-art (SOTA) baselines, especially that are LLM-based such as BLIP-2 (Li et al., 2023), SeViLA (Yu et al., 2024a) and LLaMA-VQA (Ko et al., 2023), on 5 challenging multi-choice VideoQA benchmarks: 1) TVQA (Lei et al., 2018), requireing answer questions based on video, dialogues and scenes, featuring 152,545 QA pairs from 21,793 video clips extracted from popular TV shows. 2) STAR (Wu et al., 2024a), which is designed for spatio-temporal and relational reasoning, containing 22,670 QA pairs based on 12,672 video clips. 3) DramaQA (Choi et al., 2021), which is tailored for emotional and social reasoning, featuring 16,191 QA pairs derived from 23,239 video clips. 4) VLEP (Lei et al., 2020), which focuses on predicting future events based on video and dialogues, consisting of 28,726 QA pairs from 10,000 video clips. 5) How2QA (Li et al., 2020), which is designed for instructional video comprehension, containing 46,467 QA pairs derived from 23,228 video clips.
4.2 Main Results
@@
Models Language Model # trainable params STAR DramaQA VLEP TVQA* How2QA Int. Seq. Pre. Fea. Tot. Tot. Tot. Tot. Tot. FrozenBiLM Yang et al. (2022) DeBERTa 30M - - - - - - - 57.5 86.7 MERLOT Zha et al. (2019) RoBERTa 223M - - - - - 81.4 68.4 - - SPCRL Kim et al. (2021) BERT - - - - - - 81.0 - - - AIO Wang et al. (2023a) - 110M 47.5 50.8 47.8 44.1 47.5 - - - - ATP Buch et al. (2022) CLIP - 50.6 52.9 49.4 40.6 48.4 - - - - MIST Gao et al. (2023) - - 55.6 54.2 54.2 44.5 53.9 - - - - InternVideo Wang et al. (2022b) CLIP 1.3B 62.7 65.6 54.9 51.9 58.7 - 63.9 57.2 79.0 LLaMa-VQA Ko et al. (2023) LLaMA 4.5M 66.2 67.9 57.2 52.7 65.4 84.1 71.0 70.4 - BLIP-2 Li et al. (2023) Flan-T5 432M 52.3 54.8 49.0 51.2 51.8 - 67.0 54.5 82.2 SeViLA Yu et al. (2024a) Flan-T5 216M 63.7 70.4 63.1 62.4 64.9 - 68.9 61.6 83.6 VLAP Wang et al. (2023b) Flan-T5 188M 70.0 70.4 65.9 62.2 67.1 - 69.6 63.4 83.9 BoViLA (Ours) LLaMA 4.5M 66.9 68.0 62.0 57.2 66.4 85.2 71.2 71.6 89.4
@@ 4.3 Ablation Studies
Int. Seq. Pre. Fea. Avg.
- - - 65.6 66.0 54.7 54.9 64.1
- - 66.0 66.8 59.1 54.9 65.0
✓(NGP) - 66.6 67.5 58.7 56.7 65.7
✓(NGP) - 67.0 67.7 60.0 54.8 65.9
✓(GP) 66.9 68.0 62.0 57.2 66.4
The self-questioning and answering process can easily cause ”information leakage”, where the questioner creates degenerate questions, which are meaningless and cheatingly contain target answers. Our regularization method and EDL-based filter really help solve this problem and allow the model to generate high-quality questions and filter low-quality questions to boost answers. In Figure 3, we qualitatively demonstrate the self-generated questions, the regularization term, and the EDL-estimated uncertainty through several examples. Seed Question refers to the questions labeled in the dataset, and Answer represents corresponding answers along with respective options. Degenerate Question refers to the junk questions generated by the model when regularization is not used. Self-Generated Question 1 and 2 represent high-quality questions generated by the model equipped with regularization. These questions are semantically consistent with the video and leverage the internal knowledge of the LLMs to provide diverse questions. The uncertainty is estimated by the EDL head. In these two examples, it is evident that the question quality is negatively correlated with the uncertainty. Additionally, it is worth mentioning that since we used the gumbel-softmax function with some randomness in sampling self-generated questions, different questions are generated for the same sample during each training epoch.
Furthermore, we provide comprehensive ablation studies on STAR validation set in Table 3 to verify the effectiveness of our method quantitatively. The results clearly demonstrate: 1) Only learning from regularization about questioning allows the model to achieve a 2.2% performance boost. 2) Further learning by answering these self-generated questions contributes an additional 0.7% improvement. 3) If the answerer is allowed to backpropagate gradients to the questioner, i.e. providing feedback on question quality, performance can increase by another 0.5%. 4) Moreover, applying an EDL-based filter to progressively eliminate potential junk questions that could negatively impact the model can further enhance performance by 0.9%.
Figure 3: Examples of vanilla self-generated questions (degenerate question), self-generated questions with regularization term, and the corresponding EDL-estimated uncertainty.
@@
Linear Softmax STAR
Int. Seq. Pre. Fea. Avg.
- 24.7 23.1 22.9 21.6 23.5
66.9 68.0 62.0 57.2 66.4
We also conduct an ablation study on our improved EDL method. As shown in Table 4, without our proposed linear projection to calculate the total strength (for evidence decoupling) yet simply sum it up, the model’s training will break down. This is due to the large number of classes in the vocabulary (even though most words receive very small logits), and summing directly without considering sparsity can easily lead to numerical instability. Moreover, if we do not use our proposed Softmax-based method to decouple the computation of evidence direction, the significant information loss will also cause the model to fail to converge.
4.4 More Discussion
Our proposed EDL-based filter is built on the assumption that “EDL-estimated uncertainty is able to measure the quality of self-generated questions and, is approximately negatively correlated with it”. Here, we further explore and validate this assumption. We use $\mathcal{L}_{\mathrm{v\overline{q}a}}$ and $\mathcal{L}_{\mathrm{reg}}$ as approximate measures of the quality score for self-generated questions. $\mathcal{L}_{\mathrm{v\overline{q}a}}$ represents the similarity between the self-generated question $\overline{q}$ and the seed question $q$ in terms of KL divergence, where a high $\mathcal{L}_{\mathrm{v\overline{q}a}}$ indicates a completely uncontrolled self-generated question. $\mathcal{L}_{\mathrm{reg}}$ represents the likelihood that the answerer successfully predicts the target answer from the self-generated question, where a high $\mathcal{L}_{\mathrm{reg}}$ is likely to indicates that there is no meaningful semantics in the self-generated question. $\mathcal{L}_{\mathrm{v\overline{q}a}}$ and $\mathcal{L}_{\mathrm{reg}}$ complement and reinforce each other, so we use them together to represent the quality of self-generated questions. As shown in Figure 4, both $\mathcal{L}_{\mathrm{v\overline{q}a}}$ and $\mathcal{L}_{\mathrm{reg}}$ demonstrate the pattern that ”the lower the quality of the self-generated question, the higher the uncertainty”. This insight further provides a solid explanation for the effectiveness of our proposed EDL-based filter. Moreover, we also provide a more comprehensive analysis of the uncertainty distribution modeled by our improved EDL in Appendix C, further supporting the effectiveness of the EDL-based filter.
@@ 4.4 More Discussion
5 Conclusion, Limitation and Future Work
In this paper, we introduce BoViLA, a novel video-language alignment framework that leverages self-questioning and answering to autonomously generate new training samples, thereby enhancing data utilization. Notably, this process does not rely on additional annotated data, instead fully utilizing the rich information within videos and the internal knowledge of LLMs.
A key contribution of our framework is the improvement and integration of EDL method. EDL plays a crucial role in assessing and filtering the quality of self-generated questions, ensuring that the model focuses on high-quality, informative questions. By incorporating EDL, BoViLA improves the overall alignment between video and language modalities. This integration of EDL significantly boosts the framework’s performance, allowing it to achieve superior results across five VideoQA benchmarks, outperforming current state-of-the-art methods with only a few trainable parameters.
Despite its achievements, BoViLA faces limitations due to its constrained question generation space. The questioner generates questions autoregressively based on the context of the seed question and answer, which restricts the diversity of the generated questions. Future work will focus on exploring methods to sample more diverse questions, aiming to further enhance the framework’s capabilities and expand its application scope.
@@ Appendix B Details of EDL Loss
$\mathcal{L}_{\mathrm{vqa}}^{\mathrm{edl}}=\sum_{j=1}^{N_{a}}E_{\mathrm{Dir}}[-\log p_{j{a_{j}}}]=\sum_{j=1}^{N_{a}}\int-\log p_{j{a_{j}}}\frac{1}{B(\boldsymbol{\alpha}_{j})}\prod_{i=1}^{K}p_{ji}^{\alpha_{ji}-1}d\bm{p}_{j}.$ (A)
@@
$\mathcal{L}_{\mathrm{reg}}^{\mathrm{edl}}$ $=KL[Dir(\boldsymbol{p}|\boldsymbol{\alpha})\|Dir(\boldsymbol{p}|\langle 1,\ldots,1\rangle)]$ (D)
$=E_{Dir(\boldsymbol{p}|\boldsymbol{\alpha})}\left[\log\frac{Dir(\boldsymbol{p}|\boldsymbol{\alpha})}{Dir(\boldsymbol{p}|\langle 1,\ldots,1\rangle))}\right]$ (E)
$=E_{Dir(\boldsymbol{p}|\boldsymbol{\alpha})}\left[\log Dir(\boldsymbol{p}|\boldsymbol{\alpha})-\log Dir(\boldsymbol{p}|\langle 1,\ldots,1\rangle)\right)]$ (F)
$=E_{Dir(\boldsymbol{p}|\boldsymbol{\alpha})}\left[-\log B(\boldsymbol{\alpha})+\sum_{k=1}^{K}(\alpha_{k}-1)\log p_{k}\right]+\log B(\langle 1,\ldots,1\rangle)$ (G)
$=\log\frac{B(\langle 1,\ldots,1\rangle)}{B(\boldsymbol{\alpha})}+E_{Dir(\boldsymbol{p}|\boldsymbol{\alpha})}\left[\sum_{k=1}^{K}(\alpha_{k}-1)\log p_{k}\right]$ (H)
@@ Appendix C Validity of EDL-Estimated Uncertainty
C.a Visualization of Uncertainty Distribution.